(c) Use a 0.2 to compute the exponential smoothing values for the time series. Compute MSE and a forecast for week 7. If required, round your answers to two decimal places. Week Time Series Value Forecast 18 13 16 2 18 15.5 15.75 3 40 5 6 17 13.36 14 15.19 MSE: 12.47 The forecast for week 7: 12.47

Use α = 0.2 to compute the exponential smoothing values for the time series. Compute MSE and a forecast for week 7. If required, round your answers to two decimal places. What am i doing wrong in this problem?
Answer #1

According to the given information , if 0.2, the exponential smoothing time series is determined as:

F_{t}=\alpha T_{t-1}+(1-\alpha )F_{t-1}

Forecast value of different week are as follows:

F_{2}=\alpha T_{2-1}+(1-\alpha )F_{2-1}

  0.2 × 18 + (1-02) × 18

  3.6 14.4

  18

F_{3}=\alpha T_{3-1}+(1-\alpha )F_{3-1}

  0.2 × 13 + (1-02) × 18

  2.6 + 14.4

  17

F_{4}=\alpha T_{4-1}+(1-\alpha )F_{4-1}

  =0.2\times 16+(1-0.2)\times 17

  3.2+ 13.6

  =16.8

F_{5}=\alpha T_{5-1}+(1-\alpha )F_{5-1}

  =0.2\times 11+(1-0.2)\times 16.8

  =2.2+13.44

  =15.64

F_{6}=\alpha T_{6-1}+(1-\alpha )F_{6-1}

  =0.2\times 17+(1-0.2)\times 15.64

  3.4+ 12.512

  =15.912

Therefore forecast value of 7th week is determined as:

F_{7}=\alpha T_{7-1}+(1-\alpha )F_{7-1}

  =0.2\times 14+(1-0.2)\times 15.9120

  =2.8+12.7296

  =15.5296

Week

Time Series Value

(T_{t})

Forecast

(F_{t})

Error

\left (E \right )

\left (E^{2} \right )

1

18

-

2

13

18

-5

25

3

16

17

-1

1

4

11

16.8

-5.8

33.64

5

17

15.64

1.36

1.8496

6

14

15.912

-1.912

3.655744

7

15.5296

65.14534

Now the mean square error is determined as:

MSE=\frac{\sum E^{2}}{5}=\frac{65.145344 }{5}=13.02906

Therefore the

MSE=13.0291